# A Large Language Model Evolutionary Algorithm for Automatically Generating Metaheuristics

### v3

@@ V-B Mutation Rates
There are a few interesting observations we can make: The difference between parent and offspring for GPT-3.5 is on average much smaller than for other models, and the $(1+1)$-GPT3.5-LLaMEA shows much higher differences in the beginning of the run than the other strategies with the same model. GPT-4-LLaMEA in general shows the largest differences (exploration) over the runs. It is very interesting to observe that the ratio of code differences in most cases seems to converge, indicating more exploration in the beginning of the search and more exploitation during the final parts of the search. This is interesting as the LLM has no information on the search budget $T$ of Algorithm 1, and can therefore also not base its decision to make large or small refinements on the stage of the optimization run. When we look at specific code-diffs between parents and offspring generated by the $(1+1)$-GPT-4-LLaMEA, we observe that the LLM mutates both hyperparameter values and higher level logic, like introducing a new crossover or mutation operator. In addition it mutates the comments in the code to reflect and argue about the changes. See the supplemental material for some specific examples.
@@
Figure 5 illustrates the Jaro similarity mean (five runs) of the subsequently generated algorithm names (i.e., $\text{Jaro}(\text{name}(a^{\prime}),\text{name}(a_{t+1}))$ for $t=0,\ldots,T-1$, where $a^{\prime}=a_{b}$ for $(1+1)$-selection and $a^{\prime}=a_{t}$ for $(1,1)$-selection) over iterations $t$ for each of the nine LLM-based optimization run configurations. In Figure 6, we show the Jaro similarity for just a single run selected for three of the nine configurations.
The Jaro similarity scores consistently show a behavior similar to the code difference ratios. Especially from the visualisation with individual runs it is clear that some steps only involve very small mutations (especially for GPT3.5) and sometimes a kind of restart occurs where the similarity score reaches zero.
@@
Figure 6: Jaro similarity scores between parent and offspring for each run ($5$ independent runs) of different models and strategies. Left: $(1+1)$-GPT4-ES; middle: $(1,1)$-GPT3.5-ES; right: $(1,1)$-GPT4o-ES with Details.
@@ VI Analysis of Best Algorithms
The experiment above resulted in 3 65736573\,6573 657 algorithms (out of a maximum2225555 runs with $T=100$ iterations, each, for $9$ different models. of 4 50045004\,5004 500) that were at least able to get an AOCC score larger than zero on the BBOB suite of optimization problems in $5$ dimensions. In the next step to validate our proposed framework, we evaluate how much the best of these algorithms can generalize beyond the evaluation performed during the optimization loop, and subsequently we analyze in-depth the code and behaviour of the best-performing algorithm that beats the state-the-art baselines. The evaluation is done in $5$, $10$ and $20$ dimensions, using $5$ different instances and $5$ independent runs per instance (so $25$ runs per BBOB function to optimize). For a fair comparison, we decided to re-run the generated algorithms for $d=5$ in this setting (with $25$ runs), too. The state-of-the-art baselines are the Covariance Matrix Adaptation Evolution Strategy (CMA-ES) [49], Differential Evolution (DE) [50] and an extensively optimized version of Modular CMA-ES [51, 45] denoted as CMA-best. CMA-best is a CMA-ES algorithm with active update, a Gaussian base sampler, $\lambda=20,\mu=10$, IPOP and mirrored sampling strategy and CSA step-size-adaptation (see [51] for the details of these configurations). The optimized CMA-ES algorithm is the best-configured algorithm (in terms of AOCC) on BBOB in $5$ dimensions and for a budget of 10 0001000010\,00010 000 function evaluations, out of more than 52 1285212852\,12852 128 CMA-ES and DE algorithm variants. Beating this optimized baseline is incredibly hard as it was the result of a very large modular optimization experiment. We use the CMA-best baseline to obtain a reasonable upper-bound of what is possible when optimizing an algorithm configuration to a specific benchmark set and fixed budget and dimensionality.
@@
Figure 7: The empirical attainment functions (EAF) estimate the fraction of runs that attain an arbitrary target value not later than a given runtime. The fraction ($y$-axis) denotes the cumulative fraction of target values that have been reached by an optimization algorithm, as a function of the number of function evaluations ($x$-axis). We show the EAF for the best algorithms per LLM configuration (i.e., nine algorithms) and the three baseline algorithms (CMA-ES, best CMA-ES, Differential Evolution DE), averaged over all 24 BBOB functions in $5d$.
In Figure 7 the empirical attainment functions (EAF) [47] are shown for the best algorithm generated per LLM and strategy, resulting in $9$ automatically generated algorithms plus the $3$ baselines (CMA-ES, CMA-ES-best, DE). The empirical attainment function estimates the fraction of runs that attain an arbitrary target value not later than a given runtime (i.e., number of function evaluations). Taking the partial integral of the EAF results in a more accurate version of the Empirical Cumulative Distribution Function, since it does not rely on discretization of the targets. The ”fraction” ($y$-axis in Figure 7) denotes the cumulative fraction of target values that have been reached by an optimization algorithm, as a function of the number of function evaluations ($x$-axis in Figure 7).
The area under this curve is a similar measure as the AOCC used during the LLaMEA optimization runs, and denotes how well an optimization algorithm performed over the complete run (on average over all instances, independent runs and objective functions). From the EAF curves and the area under these curves (see Table I), we can observe that LLaMEA has found one algorithm (ERADS_QuantumFluxUltraRefined, for short ERADS) with better performance for $d=5$ than the state-of-the-art CMA-ES in terms of AOCC, two algorithms (ERADS and AdaptiveDifferentialEvolutionHarmonySearch) that on average perform better than the optimized CMA-best after the total budget in $5d$ (the EAF curve reaches higher), two other algorithms that perform better than CMA-best in the first $1000$ evaluations (EnhancedFireworkAlgorithmWithLocalSearch and QuantumDifferentialParticleOptizerWithElitism) and seven algorithms that perform better than the DE baseline but worse than the CMA baseline. Detailed convergence curves per BBOB function (for $d=5$) are shown in Figure 8. From this Figure we can observe that especially for BBOB $f_{17}$ and $f_{18}$, the ERADS algorithm shows very promising search behaviour.
Figure 8: Median best-so-far function value ($y$-axis) over the 10 0001000010\,00010 000 function evaluations ($x$-axis) per BBOB function in $5d$ for the best algorithm per LLM configuration (9 algorithms: combinations of (1,1) with details, without details, and the (1+1)-variant, each with three different LLMs) and the baselines: CMA-ES, DE and CMA-best, for each of the 24 BBOB test functions.
@@
ERADS Quantum Flux Ultra Refined
@@
While the name sounds rather futuristic and very sophisticated, upon a close inspection of the code the ”ERADS Quantum Flux Ultra Refined” algorithm looks very similar a standard Differential Evolution algorithm. According to the LLM, ERADS stands for ”Enhanced RADEDM with Strategic Mutation”, and RADEDM stands for ”Refined Adaptive Differential Evolution with Dynamic Memory”. The pseudocode of ERADS, representing the key ideas of the generated Python code, was extracted manually and is shown in Algorithm 2. The full Python code can be retrieved from our online repository [39].
The main differences between ERADS and DE are the use of a memory factor to guide the mutation in certain directions and an adaptive $F$ and $CR$ strategy. It is unclear, why the LLM generated the ”Quantum Flux” component in the algorithm’s name, but it can be assumed that LLM hallucinations [52] will also occur when applied to code generation tasks.
@@
The key differences primarily revolve around the use of a memory vector and a memory factor in combination with parameter adaptation. JADE uses an optional archive with the best solutions which primarily serves to maintain diversity. Furthermore, JADE also features adaptive control parameters ($F$ and $CR$) that adjust dynamically based on the success rates of previous generations, aiming to optimise these parameters in real time to improve performance. It is interesting to observe that most of the best algorithms proposed by the LLM are similar to Differential Evolution (or Hybrids of CMA and DE), which could indicate a bias towards this kind of algorithms in the LLM, or it could indicate that overall these kind of algorithms work well for the specific benchmark suite.
VI-A Performance Analysis in Higher Dimensions
@@
The best-found algorithms using the LLaMEA framework are evaluated against the most relevant state-of-the-art optimizers, namely the previously used baselines CMA-best, CMA-ES and Differential Evolution (DE). CMA-ES and DE are using recommended hyperparameter settings. The baselines all originate from the IOHanalyzer benchmark data set [54]. In Figure 9, the empirical attainment function of the proposed algorithms and baselines are shown for dimension $d\in\{10,20\}$ using results from all BBOB functions, $5$ instances per function and $5$ random seeds ($25$ runs per BBOB function in total). It is interesting to observe that while ERADS performs well in $5$ dimensional problems, it fails to generalize well to higher dimensions, while EnhancedFireworkAlgorithmWithLocalSearch and QuantumDifferentialParticleOptizerWithElitism both have very good performance in higher dimensions for the first $2000$ function evaluations (even better than CMA-best), but around 10 0001000010\,00010 000 function evaluations CMA-ES outperforms the other algorithms clearly. It is also interesting to observe that the optimized CMA-best performs less well than a CMA-ES with default hyperparameters in $10d$ and $20d$. Just like the LLaMEA proposed algorithms, CMA-ES best was optimized on $5d$ and with a budget of 10 0001000010\,00010 000 evaluations. When looking at convergence curves per function, we note that CMA-ES outperforms mainly on the very hard functions in the benchmark ($f_{23}$ and $f_{24}$). Detailed convergence curves per BBOB function in $5d$, $10d$ and $20d$ are available in the supplemental material [39].
@@ VII Conclusions and Outlook
The LLaMEA framework proved capable of generating and evolving algorithms that perform comparably to traditional state-of-the-art metaheuristics, demonstrating the potential of LLMs to understand and innovate within the algorithmic design space effectively. Algorithms evolved through our framework successfully compete against established metaheuristics, highlighting the practical applicability of using LLMs for automated construction of optimization heuristics. Since LLMs have been trained on an exceedingly large code base, including the currently available metaheuristics, we can explain their observed strong performance by interpreting them as a universal modular framework for algorithm construction (see also Section I) that have access to a huge number of “modules” that can be combined to form new algorithms. In addition, we observed that the LLM is able to both fine-tune algorithm parameters, as well as introduce new logic such as different mutation and crossover strategies in the generated algorithms.
Considering how much work goes into the manual design of ”nature-inspired heuristics”, which are often not novel [55, 56, 57, 58] and often cannot compete with random search [10], we argue that an automated design approach for specific application domains will likely be the method of choice from now on.
@@
Supplementary material
TABLE II: The 24 BBOB noiseless functions grouped in five function categories, see [30] for details.

### v4

@@ V-B Mutation Rates
There are a few interesting observations we can make: The difference between parent and offspring for GPT-3.5 is on average smaller than for other models, and the $(1+1)$-GPT3.5-LLaMEA shows much higher differences in the beginning of the run than the other strategy with the same model. GPT-4-LLaMEA in general shows the largest differences (exploration) over the runs. It is very interesting to observe that the ratio of code differences in most cases seems to converge, indicating more exploration in the beginning of the search and more exploitation during the final parts of the search. This is interesting as the LLM has no information on the search budget $T$ of Algorithm 1, and can therefore also not base its decision to make large or small refinements on the stage of the optimization run. When we look at specific code-diffs between parents and offspring generated by the $(1+1)$-GPT-4-LLaMEA, we observe that the LLM mutates both hyper-parameter values and higher level logic, like introducing a new crossover or mutation operator. In addition it mutates the comments in the code to reflect and argue about the changes. See the supplemental material for some specific examples.
@@
Figure 6 illustrates the Jaro similarity mean (five runs) of the subsequently generated algorithm names (i.e., $\text{Jaro}(\text{name}(a^{\prime}),\text{name}(a_{t+1}))$ for $t=0,\ldots,T-1$, where $a^{\prime}=a_{b}$ for $(1+1)$-selection and $a^{\prime}=a_{t}$ for $(1,1)$-selection) over iterations $t$ for each of the LLM-based optimization run configurations.
The Jaro similarity scores consistently show a behavior similar to the code difference ratios. It is clear that some steps only involve very small mutations (especially for GPT3.5) and sometimes a kind of restart occurs where the similarity score reaches zero in a single run.
@@ VI Analysis of Best Algorithms
The experiment above resulted in 3 65736573\,6573 657 algorithms (out of a maximum2225555 runs with $T=100$ iterations, each, for $9$ different models and strategies. of 4 50045004\,5004 500) that were at least able to get an $AOCC$ score larger than zero on the BBOB suite of optimization problems in $d=5$ dimensions. In the next step to validate our proposed framework, we evaluate how much the best of these algorithms can generalize beyond the evaluation performed during the optimization loop, and subsequently we analyze in-depth the code and behaviour of the best-performing algorithm that beats the state-the-art baselines. The evaluation is done in $5$, $10$ and $20$ dimensions, using $5$ different instances and $5$ independent runs per instance (so $25$ runs per BBOB function to optimize). For a fair comparison, we decided to re-run the generated algorithms for $d=5$ in this setting (with $25$ runs), too. The state-of-the-art baselines are the Covariance Matrix Adaptation Evolution Strategy (CMA-ES) [49], Differential Evolution (DE) [50] and an extensively optimized version of Modular CMA-ES [51, 46] denoted as CMA-best. CMA-best is a CMA-ES algorithm with active update, a Gaussian base sampler, $\lambda=20,\mu=10$, IPOP and mirrored sampling strategy and CSA step-size-adaptation (see [51] for the details of these configurations). The optimized CMA-ES algorithm is the best-configured algorithm (in terms of $AOCC$) on BBOB in $5$ dimensions and for a budget of 10 0001000010\,00010 000 function evaluations, out of more than 52 1285212852\,12852 128 CMA-ES and DE algorithm variants. Beating this optimized baseline is incredibly hard as it was the result of a very large modular optimization experiment. We use the CMA-best baseline to obtain a reasonable upper-bound of what is possible when optimizing an algorithm configuration to a specific benchmark set and fixed budget and dimensionality.
@@
Figure 7: The empirical attainment functions (EAF) estimate the fraction of runs that attain an arbitrary target value not later than a given runtime. The fraction ($y$-axis) denotes the cumulative fraction of target values that have been reached by an optimization algorithm, as a function of the number of function evaluations ($x$-axis). We show the EAF for the best algorithms per LLM configuration (i.e., six algorithms) and the three baseline algorithms (CMA-ES, best CMA-ES, Differential Evolution (DE)), averaged over all 24 BBOB functions in $5d$.
In Figure 7 the empirical attainment functions (EAF) [24] are shown for the best algorithm generated per LLM and strategy, resulting in $6$ automatically generated algorithms plus the $3$ baselines (CMA-ES, CMA-ES-best, DE). The empirical attainment function estimates the fraction of runs that attain an arbitrary target value not later than a given runtime (i.e., number of function evaluations). Taking the partial integral of the EAF results in a more accurate version of the Empirical Cumulative Distribution Function, since it does not rely on discretization of the targets. The ”fraction” ($y$-axis in Figure 7) denotes the cumulative fraction of target values that have been reached by an optimization algorithm, as a function of the number of function evaluations ($x$-axis in Figure 7).
Such EAF curves summarize how well an optimization algorithm performs over the complete run for the given set of target values, again on average over all instances, independent runs and objective functions. From the EAF curves and the area under these curves (see Table I), we can observe that LLaMEA has found one algorithm (ERADS_QuantumFluxUltraRefined, for short ERADS) with better performance for $d=5$ than the state-of-the-art CMA-ES in terms of $AOCC$, two algorithms (ERADS and AdaptiveDifferentialEvolutionHarmonySearch) that on average perform better than the optimized CMA-best after the total budget in $5d$ (the EAF curve reaches higher), two other algorithms that perform better than CMA-best in the first $1000$ evaluations (EnhancedFireworkAlgorithmWithLocalSearch and QuantumDifferentialParticleOptizerWithElitism) and seven algorithms that perform better than the DE baseline but worse than the CMA baseline. Detailed convergence curves per BBOB function (for $d=5$) for ERADS_QuantumFluxUltraRefined and three baselines are shown in Figure 8, these curves are also available for all best algorithms in the supplemental material. From this Figure we can observe that especially for BBOB $f_{17}$ and $f_{18}$, the ERADS algorithm shows very promising search behaviour.
Refer to captionRefer to caption
Figure 8: Median best-so-far function value ($y$-axis) over the 10 0001000010\,00010 000 function evaluations ($x$-axis) per BBOB function in $5d$ for the ERADS_QuantumFluxUltraRefined algorithm and the baselines: CMA-ES, DE and CMA-best, for each of the 24 BBOB test functions.
@@
VI-A The ”ERADS_QuantumFluxUltraRefined” Algorithm
@@
While the name sounds rather futuristic and very sophisticated, upon a close inspection of the code the ”ERADS_QuantumFluxUltraRefined” algorithm looks very similar a standard Differential Evolution algorithm. According to the LLM, ERADS stands for ”Enhanced RADEDM with Strategic Mutation”, and RADEDM stands for ”Refined Adaptive Differential Evolution with Dynamic Memory”. The pseudocode of ”ERADS_QuantumFluxUltraRefined”, representing the key ideas of the generated Python code, was extracted manually and is shown in Algorithm 2. The full Python code can be retrieved from our online repository [40].
The main differences between ERADS and DE are the use of a memory factor to guide the mutation in certain directions and an adaptive $F$ and $CR$ control strategy. It is unclear, why the LLM generated the ”Quantum Flux” component in the algorithm’s name, but it can be assumed that LLM hallucinations [52] will also occur when applied to code generation tasks.
@@
The key differences primarily revolve around the use of a memory vector and a memory factor (lines 4, 10, 18 and 31 of Algorithm 2) in combination with parameter adaptation (line 13 of Algorithm 2). Moreover, ERADS involves three randomly selected individuals from the population, the best individual, and the memory vector in the mutation operator (line 18) and then, with probability $CR=0.95$, performs a crossover of the mutant and the actual individual (line 20). This crossover operator is the classical binomial crossover from DE, which copies a variable $\mathbf{v}_{i,j}$ from the mutant vector $\mathbf{v}_{i}$ with probability $CR=0.95$, i.e., the trial vector in line 18 is almost always identical to the mutant vector. In classical DE notation, the mutant $\mathbf{v}_{i}$ ($i$ being the index of individuals in the population) is generated in Algorithm 2 as follows:
$\mathbf{v}_{i}=\mathbf{x}_{r_{1}}+F_{g}\cdot(\mathbf{x}_{best}-\mathbf{x}_{r_{1}})+F_{g}\cdot(\mathbf{x}_{r_{2}}-\mathbf{x}_{r_{3}})+F_{g}\cdot m_{f}\cdot\mathbf{m}$
As usual, the $r_{i}$ indicate randomly sampled solution indices from the current population and $F_{g}$ is the mutation factor ($F_{current}$ in Algorithm 2), which in ERADS changes per generation $g$. The first two terms and crossover can best be characterized as a DE/rand-to-best/1/bin strategy in classical DE notation [54], but the additional memory vector that contributes $0.3\cdot F_{g}\cdot\mathbf{m}$ (remember that $m_{f}=0.3$, as initialized in line 4, remains constant) is new. It should also be noted that ERADS updates $\mathbf{x}_{best}$ immediately within the current generation, if the newly generated individual improves $\mathbf{x}_{best}$. Similarly, $\mathbf{m}$ is updated immediately, if the newly generated individual improves $\mathbf{x}_{best}$, as follows:
$\mathbf{m}\leftarrow 0.7\cdot\mathbf{m}+0.3\cdot F_{g}\cdot(\mathbf{v}_{i}-\mathbf{x}_{i})$
It is interesting to observe that this update does not use the trial vector that was generated by crossover of $\mathbf{v}_{i}$ and $\mathbf{x}_{i}$ in line 20, and which is used to update $\mathbf{x}_{best}$ upon an improvement. Instead, it just uses the mutant vector $\mathbf{v}_{i}$, which, however, is almost identical to the trial vector due to $CR=0.95$, which is the probability for each component of the mutant vector to be copied to the trial vector.
Finally, ERADS also includes a generational adaptation that linearly increases $F_{g}$ from an initial value of $0.55$ to the final value of $0.85$, i.e., $F_{g}=0.55+0.3\cdot t/B$ (line 13). While this increases the mutation factor over time, it still remains below the maximum range of the mutation factor that is typically recommended, i.e., $[0,1]$ [53].
To further examine the hyper-parameters of the ERADS algorithm we include additional hyper-parameter optimization experiments in the supplemental material. The result clarifies that the LLM has already generated well performing hyper-parameter settings for $d=5$ that can not be improved further.
@@
In contrast to ERADS, JADE uses an optional archive with the best solutions which primarily serves to maintain diversity. Furthermore, JADE also features adaptive control parameters ($F$ and $CR$) that adjust dynamically based on the success rates of previous generations, aiming to optimise these parameters in real time to improve performance.
It is interesting to observe that most of the best algorithms proposed by the LLM are similar to Differential Evolution (or hybrids of CMA and DE), which could indicate a bias towards this kind of algorithms in the LLM, or it could indicate that overall this kind of algorithms work well for the specific benchmark suite.
VI-B Performance Analysis in Higher Dimensions
@@
The best-found algorithms using the LLaMEA framework are evaluated against the most relevant state-of-the-art optimizers, namely the previously used baselines CMA-best, CMA-ES and Differential Evolution (DE). CMA-ES and DE are using recommended hyper-parameter settings. The baselines all originate from the IOHanalyzer benchmark data set [55]. In Figure 10, the empirical attainment function of the proposed algorithms and baselines are shown for dimension $d\in\{10,20\}$ using results from all BBOB functions, $5$ instances per function and $5$ random seeds ($25$ runs per BBOB function in total). It is interesting to observe that while ERADS performs well in $5$ dimensional problems, it fails to generalize well to higher dimensions, while EnhancedFireworkAlgorithmWithLocalSearch and QuantumDifferentialParticleOptizerWithElitism both have very good performance in higher dimensions for the first $2000$ function evaluations (even better than CMA-best), but around 10 0001000010\,00010 000 function evaluations CMA-ES outperforms the other algorithms clearly. It is also interesting to observe that the optimized (for $d=5$) CMA-best performs less well than a CMA-ES with default hyper-parameters in $10d$ and $20d$. Just like the LLaMEA proposed algorithms, CMA-ES best was optimized on $5d$ and with a budget of 10 0001000010\,00010 000 evaluations. When looking at convergence curves per function, we note that CMA-ES outperforms mainly on the very hard functions in the benchmark ($f_{23}$ and $f_{24}$). Detailed convergence curves per BBOB function in $5d$, $10d$ and $20d$ are available in the supplemental material [40].
@@ VII Conclusions and Outlook
The LLaMEA framework proved capable of generating and evolving algorithms that perform comparably to traditional state-of-the-art metaheuristics, demonstrating the potential of LLMs to innovate within the algorithmic design space effectively. Algorithms evolved through our framework successfully compete against established metaheuristics, highlighting the practical applicability of using LLMs for automated construction of optimization heuristics. Since LLMs have been trained on an exceedingly large code base, including the currently available metaheuristics, we can explain their observed strong performance by interpreting them as a universal modular framework for algorithm construction (see also Section I) that have access to a huge number of “modules” that can be combined to form new algorithms. In addition, we observed that the LLM is able to both fine-tune algorithm parameters, as well as introduce new logic such as different mutation and crossover strategies in the generated algorithms.
Considering how much work goes into the manual design of “nature-inspired heuristics”, which are often not novel [56, 57, 58, 59] and often cannot compete with random search [10], we argue that an automated design approach for specific application domains will likely be the method of choice from now on.
@@
Appendix A BBOB Function Groups
In the experiments we heavily use the black-box optimization benchmarking suite (BBOB), which consists of $24$ different functions categorized in $5$ different function groups. In Table II, each function, its name and its respective function group is given. For additional details see [61]. As it is common in BBOB, rotation and translation are applied on each of these functions to generate additional function instances. Using this wide range of different function landscapes and different random instances allows for developing (and, in this case, automatically discover) strong black-box optimization metaheuristics that perform well on a large range of problems.
TABLE II: The 24 BBOB noiseless functions grouped in five function categories
@@ Appendix A BBOB Function Groups
Appendix B Extended Results on BBOB
In the following supplementary section we provided additional results of the BBOB experiments done with the best algorithms discovered by each LLM and strategy combination. This includes the strategy $1,1$ with details. In this strategy we give the LLM additional information in the form of AOCC scores per BBOB function group, instead of one average AOCC score. Since this additional feedback did not result in better optimization performance, we only show these results in the supplemental section. In Figure 10 we can observe the performance of the $9$ algorithms that were the results of each LLaMEA configuration (three LLM versions and the three strategies (1+1, 1,1 and 1,1-with-details) against the three baselines, CMA-ES, optimized CMA-ES and DE (in $5d$, $10d$, and $20d$). In Figures 11, 12 and 13, the convergence curves (function values) per BBOB function of these algorithms are shown in respectively $5d$, $10d$ and $20d$.
Refer to captionRefer to caption
Figure 10: The empirical attainment function (EAF) estimates the percentage of runs that attain an arbitrary target value not later than a given runtime. EAF for the best algorithm per configuration and the baselines; CMA-ES, CMA-best and DE, averaged over all 24 BBOB functions in $5d$ (top plot), $10d$ (left bottom plot) and $20d$ (right bottom plot), respectively.